OpenStudy (anonymous):

If g(x) = 2x- x +5 , then g (2) =

hero (hero):

g(x) = 2x - x + 5 g(x) = x + 5 g(2) = 2 + 5 g(2) = ?

OpenStudy (anonymous):

am i supposed to add 5+2

hero (hero):

Hint: yes

OpenStudy (anonymous):

so is the ans 7